# Fusions of One-Variable First-Order Modal Logics

### 04 Mar 2026 v1

@@ 1 Introduction
The aim of this contribution is to initiate a systematic investigation into the extent to which these preservation results for fusions of propositional modal logics can be lifted to (fragments of) first-order modal logics and corresponding many-dimensional modal logics. We focus in this paper on one-variable first-order modal logics, either without equality or with equality and non-rigid constants, and on corresponding commutators and products of propositional modal logics. These logics have been studied extensively [11, 18, 24, 5] as they are often computationally and semantically better behaved than full first-order modal logics. For fusions without equality our main result is as follows:
@@ 2 Preliminaries
We consider first-order (multi-)modal logics in languages with a single, fixed, individual variable. A language $\mathcal{L}_{M}$ contains countably many unary predicate letters, the individual variable $x$, the Booleans $\neg$ and $\land$, the quantifier symbol $\exists$, and a finite set $M$ of unary (necessity-like) modal operators, denoted by $\Box$, possibly with subscripts. Thus, in $\mathcal{L}_{M}$, atomic formulas have the form $P(x)$, where $P$ is a predicate letter. Formulas of $\mathcal{L}_{M}$, denoted by $\varphi,\psi,\ldots$, are built from atomic ones in the usual way, the only allowed quantifier being $\exists x$. The symbols $\bot$, $\to$, $\vee$, $\forall$, and $\Diamond$ are standard abbreviations. We also denote $\Box^{\leq n}\varphi=\varphi\land\Box\varphi\land\dots\land\Box^{n}\varphi$. Proposition letters $p,q,\dots$ are abbreviations for $\forall x\,P(x)$, $\forall x\,Q(x)$, etc. We also consider languages with equality: $\mathcal{L}^{=}_{M}$ contains, in addition to the symbols of $\mathcal{L}_{M}$, the binary predicate symbol == and constant symbols, denoted by $c_{1},c_{2},\dots$ Thus, in $\mathcal{L}^{=}_{M}$, a term $t$ is either $x$ or a constant $c_{i}$, and atomic formulas have the form $P(t)$ or $t_{1}=t_{2}$. We usually identify languages with sets of their formulas. The set of subformulas of a formula $\varphi$ is denoted by $\mathop{\mathit{sub}}\varphi$.
By a one-variable (first-order) $M$-logic, we mean a set of $\mathcal{L}_{M}$-formulas that includes the one-variable fragment of the classical first-order logic $\mathbf{FO}$ and formulas $\Box\bigl(P(x)\to Q(x)\bigr)\to\bigl(\Box P(x)\to\Box Q(x)\bigr)$, for all $\Box\in M$, and is closed under Modus Ponens (MP), Generalisation ($\varphi/\forall x\,\varphi$), Necessitation ($\varphi/\Box\varphi$, for each $\Box\in M$), and Substitution (of arbirary one-variable formulas for formulas of the form P(x)P(x)). [^1] One-variable fragments of first-order modal logics are one-variable logics in our sense, but not every one-variable logic is a one-variable fragment of a first-order modal logic; see, e.g., [28, Section 3]. If $L$ is a one-variable $M$-logic and $\Gamma\subseteq\mathcal{L}_{M}$, then $L+\Gamma$ denotes the smallest one-variable $M$-logic including $L\cup\Gamma$; we write $L+\varphi$ for $L+\{\varphi\}$. We note that every one-variable $M$-logic contains the converse Barcan formula $\textit{CBF}_{\Box}=\Box\forall x\,P(x)\to\forall x\,\Box P(x)$, for each $\Box\in M$.
@@
In languages with equality, additionally,
@@
The fusion of Kripke frames $\mathfrak{F}_{1}=(W,\{R_{\Box}\}_{\Box\in M_{1}})$ and $\mathfrak{F}_{2}=(W,\{R_{\Box}\}_{\Box\in M_{2}})$, where $M_{1}\cap M_{2}=\varnothing$, is the frame $\mathfrak{F}_{1}\otimes\mathfrak{F}_{2}=(W,\{R_{\Box}\}_{\Box\in M_{1}\cup M_{2}})$. If $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ are classes of Kripke frames, then $\mathscr{C}_{1}\otimes\mathscr{C}_{2}$ denotes the class of frames comprising $\mathfrak{F}_{1}\otimes\mathfrak{F}_{2}$, for all $\mathfrak{F}_{1}\in\mathscr{C}_{1}$ and $\mathfrak{F}_{2}\in\mathscr{C}_{2}$ with the same set of worlds.
@@
The fusion of a propositional $M_{1}$-logic $L_{1}$ and a propositional $M_{2}$-logic $L_{2}$, where $M_{1}\cap M_{2}=\varnothing$, is smallest propositional $(M_{1}\cup M_{2})$-logic $L_{1}\otimes L_{2}$ that includes $L_{1}\cup L_{2}$.
The semicommutator [^2] In general, semicommutators are also required to contain confluence axioms for commuting modalities; these are derivable in semicommutators with $\mathbf{S5}$. and the commutator of a propositional $M$-logic $L$ with the propositional monomodal logic $\mathbf{S5}$ are the propositional $(M\cup\{\Box_{E}\})$-logics
@@ 3 Fusions of Equality-Free One-Variable Logics
Let $\Phi$ be a set of $\mathcal{L}$-formulas closed under subformulas. A $\Phi$-type $t\subseteq\{\psi,\neg\psi\mid\psi\in\Phi\}$ is such that
$\neg\psi\in t\quad$ $\text{ iff }\quad\psi\notin t,$ $\text{for each }\psi\in\Phi,$ (t¬)
$\psi_{1}\land\psi_{2}\in t\quad$ $\text{ iff }\quad\psi_{1},\psi_{2}\in t,$ $\text{for each }\psi_{1}\land\psi_{2}\in\Phi.$ (t)
The number of $\Phi$-types is bounded by $2^{|\Phi|}$ (for $\neg\psi\in\Phi$, a type contains either $\neg\psi$ or both $\neg\neg\psi$ and $\psi$). A $\Phi$-quasistate is a non-empty set $\boldsymbol{q}$ of $\Phi$-types such that, for all $t\in\boldsymbol{q}$,
@@ 3.1 Global Consequence
Define a model $\mathfrak{M}_{0}=(\mathfrak{F},\langle\mathbb{N},I_{w}\rangle_{w\in W})$ by setting $I_{w}(Q)=\{1,\ldots,k\}$ iff $w\in W^{k}$, for each $k\in\mathbb{N}$, and $I_{w}(p)=1$ iff $w\in\bigcup\{W^{i}\mid i\in\mathbb{N}\}$. Then $\mathfrak{M}_{0}\models\varphi$ and so $\varphi\not\vdash_{L_{1}\otimes L_{2}}^{\ast}\bot$. On the other hand, it should be clear that $\mathfrak{M}\not\models\varphi$, for any finite model $\mathfrak{M}$. Hence, $L_{1}\otimes L_{2}$ does not have the global $d$-fmp. ∎
@@ 3.2 Local Consequence
Let $(W,R_{i})\in\mathcal{C}_{i}$. For $w,w^{\prime}\in W$, let $\textit{dist}_{R_{i}}(w,w^{\prime})$ be the length of the shortest $R_{i}$-path from $w$ to $w^{\prime}$. A tapered $\Phi$-run through $((W,R_{i}),w_{0})$ is a function from an $R_{i}$-closed subset of $W$, denoted by $\mathop{\mathsf{dom}}r$, such that each $r(w)$ is a $\Gamma_{h(w)}^{\mathbf{i}}(\Phi)$-type, for $w\in\mathop{\mathsf{dom}}r$, where $h(w)=\textit{md}^{\mathbf{i}}(\Phi)-\textit{dist}_{R_{i}}(w_{0},w)$. For $w\in\mathop{\mathsf{dom}}r$, a run $r$ is called
@@
We inductively define a tapered $1$-cactus $\mathfrak{C}$ for $\Phi$ as a tuple of the form $((W,R_{1},R_{2}),w_{0},\mathfrak{R})$, where $(W,R_{1})$ is a $\mathcal{C}_{1}$-frame, $(W\setminus\{w_{0}\},R_{2})$ is a $\mathcal{C}_{2}$-frame, a distinguished world $w_{0}\in W$ and $\mathfrak{R}$ is a set of tapered runs that are $1$- and $2$-coherent as well as $1$-saturated on $W$ but $2$-saturated only on $W\setminus\{w_{0}\}$. A tapered $2$-cactus is defined symmetrically.
@@
A tapered $1$-cactus $((W,R_{1},R_{2}),w_{0},\mathfrak{R})$ for $\emptyset$ is constructed by first taking sufficiently large disjoint unions $(W,R_{1})$ and $(W^{\prime},R_{2}^{\prime})$ of copies of $\mathfrak{F}_{1}\in\mathcal{C}_{1}$ and $\mathfrak{F}_{2}\in\mathcal{C}_{2}$, respectively, such that $|W|-1=|W^{\prime}|$. We then take any $w_{0}\in W$ and obtain the required $(W,R_{1},R_{2})$ by applying to $R_{2}^{\prime}$ the bijection between $W\setminus\{w_{0}\}$ and $W^{\prime}$. Finally, $\mathfrak{R}$ consists of a single tapered $\emptyset$-run $r$ with $r(w)=\emptyset$, for all $w\in W$, which is trivially $1$- and $2$-coherent as well as $1$- and $2$-saturated.
@@
For each $v\in W^{0}$ with $h(v)\geq 0$, we consider the $\Theta^{1}(\Phi)$-quasistate $\boldsymbol{q}^{v}\in\boldsymbol{Q}_{1}(\varphi)$ such that $\mathfrak{M}^{0},v\models(\hat{\boldsymbol{q}}^{v})^{\mathbf{1}}$. Observe that $\boldsymbol{q}^{v}$ may contain formulas $\chi$ with $\textit{md}^{\mathbf{1}}{(\chi)}>h(v)$. Let $\Phi^{v}=\Theta^{\mathbf{1}}(\Phi)\cap\Gamma^{\mathbf{1}}_{h(v)}(\Phi)$ and consider $\boldsymbol{q}^{v}|_{\smash{\Gamma^{\mathbf{1}}_{h(v)}(\Phi)}}$, which is a $\Phi^{v}$-quasistate with $\boldsymbol{q}^{v}|_{\smash{\Gamma^{\mathbf{1}}_{h(v)}(\Phi)}}=\mathfrak{R}^{0}(v)|_{\Theta^{\mathbf{1}}(\Phi)}$. Let $\varphi^{v}$ be its realisability sentence. Since $\textit{adp}^{\mathbf{2}}(\varphi^{v})=k-1$, we apply the induction hypothesis to the $\mathcal{L}$-formula $\varphi^{v}$ and obtain a tapered 2-cactus $\mathfrak{C}^{v}=((W^{v},R_{1}^{v},R_{2}^{v}),v,\mathfrak{R}^{v})$ for $\Phi^{v}$ (strictly speaking, for $\mathop{\textit{sub}}\varphi^{v}$ but the only difference is that $\mathop{\textit{sub}}\varphi^{v}$ contains Boolean combinations of quantified formulas from $\Phi^{v}$) such that $\boldsymbol{q}^{v}|_{\smash{\Gamma^{\mathbf{1}}_{h(v)}(\Phi)}}=\mathfrak{R}^{v}(v)$. For each $v\in W^{0}$ with $h(v)<0$, we take a tapered 2-cactus $\mathfrak{C}^{v}$ for $\emptyset$. Next, for each $v\in W^{0}\setminus\{w_{0}\}$, we graft the tapered 2-cactus $\mathfrak{C}^{v}$ onto $v$ in the 1-quasimodel $\mathfrak{Q}$ to obtain a tapered 1-cactus $\mathfrak{C}$. A fully grafted tapered 1-cactus can be obtained by grafting, in addition, $\mathfrak{C}^{w_{0}}$ onto $w_{0}$ in $\mathfrak{C}$.
@@ 4 Fusions of One-Variable Logics with Equality
We show that once equality is allowed in one‑variable modal logics, the transfer results established in Section 3 no longer hold. Remarkably, the counterexamples we introduce arise quite naturally. The propositional modal logic $\mathbf{Diff}$ of the ‘elsewhere’ operator was introduced by Von Wright [31] and is determined by the class $\mathcal{D}$ of all disjoint unions of frames $(W,R)$ with $R=\{(w,v)\in W^{2}\mid w\not=v\}$; see also [8]. We are interested here in the one-variable modal logics $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}$ and $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\text{fin}}$ with $\mathcal{D}_{\text{fin}}$ the class of disjoint unions of finite frames in $\mathcal{D}$. Note that the propositional logic Diff has the finite model property, and so is complete for both $\mathcal{D}$ and $\mathcal{D}_{\text{fin}}$. On the other hand, while one can easily prove the finite model property for $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\text{fin}}$, the logic $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}$ does not have the finite model property, and so $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}\subsetneq\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\text{fin}}$ [20]. It has been observed [19, 20] that both logics can be reduced in polytime to the two-variable fragment of first-order logic with counting on arbitrary and finite domains, respectively, and so they are decidable in coNExpTime.
@@
If $e$ if the form $y=n$, then we take $n$ constants $c_{1}^{e},\ldots,c_{n}^{e}$ interpreted differently and make the set of their interpretations the extension of $P_{y}$:
@@
Figure 1: Proof of Theorem 4.1. a) Formula $\textsf{card}(I,P)$. b) Equation $e$ of the form $y=z_{1}\times z_{2}$ with $z_{1}=3$ and $z_{2}=2$ (only worlds in $W$ and their domains are shown).
@@
which implies that equation $e$ holds in the sense that $|P_{y}|_{w}=|P_{z_{1}}|_{w}\times|P_{z_{2}}|_{w}$, for all $w\in W$; see Fig. 1 b).

### 06 Mar 2026 v2

@@ 1 Introduction
The aim of this contribution is to initiate a systematic investigation into the extent to which these preservation results for fusions of propositional modal logics can be lifted to (fragments of) first-order modal logics and corresponding many-dimensional modal logics. We focus in this paper on one-variable first-order modal logics, either without equality or with equality and non-rigid constants, and on corresponding commutators and products of propositional modal logics. These logics have been studied extensively [11, 18, 24, 5, 28] as they are often computationally and semantically better behaved than full first-order modal logics. For fusions without equality our main result is as follows:
@@ 2 Preliminaries
We consider first-order (multi-)modal logics in languages with a single, fixed, individual variable. A language $\mathcal{L}_{M}$ contains countably many unary predicate letters, the individual variable $x$, the Booleans $\neg$ and $\land$, the quantifier symbol $\exists$, and a finite set $M$ of unary (necessity-like) modal operators, denoted by $\Box$, possibly with subscripts. Thus, in $\mathcal{L}_{M}$, atomic formulas have the form $P(x)$, where $P$ is a predicate letter. Formulas of $\mathcal{L}_{M}$, denoted by $\varphi,\psi,\ldots$, are built from atomic ones in the usual way, the only allowed quantifier being $\exists x$. The symbols $\bot$, $\to$, $\vee$, $\forall$, and $\Diamond$ are standard abbreviations. We also denote $\Box^{\leq n}\varphi=\varphi\land\Box\varphi\land\dots\land\Box^{n}\varphi$. Proposition letters $p,q,\dots$ are abbreviations for $\exists x\,P(x)$, $\exists x\,Q(x)$, etc. We also consider languages with equality: $\mathcal{L}^{=}_{M}$ contains, in addition to the symbols of $\mathcal{L}_{M}$, the binary predicate symbol == and constant symbols, denoted by $c_{1},c_{2},\dots$ Thus, in $\mathcal{L}^{=}_{M}$, a term $t$ is either $x$ or a constant $c_{i}$, and atomic formulas have the form $P(t)$ or $t_{1}=t_{2}$. We usually identify languages with sets of their formulas. The set of subformulas of a formula $\varphi$ is denoted by $\mathop{\mathit{sub}}\varphi$.
By a one-variable (first-order) $M$-logic, we mean a set of $\mathcal{L}_{M}$-formulas that includes the one-variable fragment of the classical first-order logic $\mathbf{FO}$ and formulas $\Box\bigl(P(x)\to Q(x)\bigr)\to\bigl(\Box P(x)\to\Box Q(x)\bigr)$, for all $\Box\in M$, and is closed under Modus Ponens (MP), Generalisation ($\varphi/\forall x\,\varphi$), Necessitation ($\varphi/\Box\varphi$, for each $\Box\in M$), and Substitution (of arbirary one-variable formulas for formulas of the form P(x)P(x)). [^1] One-variable fragments of first-order modal logics are one-variable logics in our sense, but not every one-variable logic is a one-variable fragment of a first-order modal logic; see, e.g., [29, Section 3]. If $L$ is a one-variable $M$-logic and $\Gamma\subseteq\mathcal{L}_{M}$, then $L+\Gamma$ denotes the smallest one-variable $M$-logic including $L\cup\Gamma$; we write $L+\varphi$ for $L+\{\varphi\}$. We note that every one-variable $M$-logic contains the converse Barcan formula $\textit{CBF}_{\Box}=\Box\forall x\,P(x)\to\forall x\,\Box P(x)$, for each $\Box\in M$.
@@
In languages $\mathcal{L}_{M}^{=}$ with equality, additionally,
@@
The fusion of Kripke frames $\mathfrak{F}_{1}=(W,\{R_{\Box}\}_{\Box\in M_{1}})$ and $\mathfrak{F}_{2}=(W,\{R_{\Box}\}_{\Box\in M_{2}})$, where $M_{1}\cap M_{2}=\emptyset$, is the frame $\mathfrak{F}_{1}\otimes\mathfrak{F}_{2}=(W,\{R_{\Box}\}_{\Box\in M_{1}\cup M_{2}})$. If $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ are classes of Kripke frames, then $\mathscr{C}_{1}\otimes\mathscr{C}_{2}$ denotes the class of frames comprising $\mathfrak{F}_{1}\otimes\mathfrak{F}_{2}$, for all $\mathfrak{F}_{1}\in\mathscr{C}_{1}$ and $\mathfrak{F}_{2}\in\mathscr{C}_{2}$ with the same set of worlds.
@@
The fusion of a propositional $M_{1}$-logic $L_{1}$ and a propositional $M_{2}$-logic $L_{2}$, where $M_{1}\cap M_{2}=\emptyset$, is smallest propositional $(M_{1}\cup M_{2})$-logic $L_{1}\otimes L_{2}$ that includes $L_{1}\cup L_{2}$.
The semicommutator [^2] In general, semicommutators are also required to contain confluence axioms for commuting modalities; these are derivable in semicommutators with $\mathbf{S5}$. and the commutator  [23, 25] of a propositional $M$-logic $L$ with the propositional monomodal logic $\mathbf{S5}$ are the propositional $(M\cup\{\Box_{E}\})$-logics
@@ 3 Fusions of Equality-Free One-Variable Logics
Let $\Phi$ be a finite set of $\mathcal{L}$-formulas closed under subformulas. A $\Phi$-type $t$ is a maximal Boolean-consistent subset of $\{\psi,\neg\psi\mid\psi\in\Phi\}$, that is,
$\neg\psi\in t\quad$ $\text{ iff }\quad\psi\notin t,$ $\text{for each }\psi\in\Phi,$ (t1)
$\psi_{1}\land\psi_{2}\in t\quad$ $\text{ iff }\quad\psi_{1},\psi_{2}\in t,$ $\text{for each }\psi_{1}\land\psi_{2}\in\Phi.$ (t2)
The number of $\Phi$-types is bounded by $2^{|\Phi|}$ (for $\neg\psi\in\Phi$, a type contains either $\neg\psi$ or both $\neg\neg\psi$ and $\psi$). A $\Phi$-quasistate is a non-empty set $\boldsymbol{q}$ of $\Phi$-types such that
@@ 3.1 Global Consequence
Define a model $\mathfrak{M}_{0}=(\mathfrak{F},\langle\mathbb{N},I_{w}\rangle_{w\in W})$ by setting $I_{w}(Q)=\{1,\ldots,k\}$ iff $w\in W^{k}$, for each $k\in\mathbb{N}$, and $I_{w}(p)=1$ iff $w\in\bigcup_{i\in\mathbb{N}}W^{i}$. Then $\mathfrak{M}_{0}\models\varphi$ and so $\varphi\not\vdash_{L_{1}\otimes L_{2}}^{\ast}\bot$. On the other hand, it should be clear that $\mathfrak{M}\not\models\varphi$, for any finite model $\mathfrak{M}$. Hence, $L_{1}\otimes L_{2}$ does not have the global $d$-fmp. ∎
@@ 3.2 Local Consequence
Let $(W,R_{i})\in\mathcal{C}_{i}$. For $w,w^{\prime}\in W$, let $\textit{dist}_{R_{i}}(w,w^{\prime})$ be the length of the shortest $R_{i}$-path from $w$ to $w^{\prime}$. A tapered $\Phi$-run through $((W,R_{i}),w_{0})$ is a function $r$ from an $R_{i}$-closed subset of $W$, denoted by $\mathop{\mathsf{dom}}r$, such that each $r(w)$ is a $\Gamma_{h(w)}^{\mathbf{i}}(\Phi)$-type, where $h(w)=\textit{md}^{\mathbf{i}}(\Phi)-\textit{dist}_{R_{i}}(w_{0},w)$. For $w\in\mathop{\mathsf{dom}}r$, run $r$ is
@@
We inductively define a tapered $1$-cactus $\mathfrak{C}$ for $\Phi$ as a tuple of the form $((W,R_{1},R_{2}),w_{0},\mathfrak{R})$, where $(W,R_{1})\in\mathcal{C}_{1}$ and $(W\setminus\{w_{0}\},R_{2})\in\mathcal{C}_{2}$ with a distinguished world $w_{0}\in W$, which is not in the domain or range of $R_{2}$, and $\mathfrak{R}$ is a set of tapered runs that are $1$- and $2$-coherent as well as $1$-saturated on $W$ but $2$-saturated only on $W\setminus\{w_{0}\}$. A tapered $2$-cactus is defined symmetrically. See Fig. 1, where we show the structure of the frame.
[…]
degenerate
2-cacti
degenerate
1-cacti
$\textit{adp}(\varphi^{w_{3}})=1$
$\textit{md}^{\mathbf{1}}(\varphi^{w_{3}})=1$
$\Phi^{w_{3}}=\$
$\Theta^{\mathbf{2}}(\Phi^{w_{0}})\cap\Gamma^{\mathbf{2}}_{0}(\Phi^{w_{0}})=\$
$\{p,\Box_{1}\neg r,\dots\}$
Figure 1: A fully grafted tapered $1$-cactus satisfying $\varphi=\Diamond_{1}\Diamond_{2}p\land\Diamond_{2}\Diamond_{2}\Diamond_{1}r$: solid arrows with dark background for $R_{1}$-components and dotted arrows with light background for $R_{2}$-components.
@@
A tapered ‘degenerate’ $1$-cactus $((W,R_{1},R_{2}),w_{0},\mathfrak{R})$ for $\emptyset$ is constructed by first taking sufficiently large disjoint unions $(W,R_{1})$ and $(W^{\prime},R_{2}^{\prime})$ of copies of $\mathfrak{F}_{1}\in\mathcal{C}_{1}$ and $\mathfrak{F}_{2}\in\mathcal{C}_{2}$, respectively, such that $|W|-1=|W^{\prime}|$. We then take any $w_{0}\in W$ and obtain the required $(W,R_{1},R_{2})$ by applying to $R_{2}^{\prime}$ the bijection between $W\setminus\{w_{0}\}$ and $W^{\prime}$. Finally, $\mathfrak{R}$ consists of a single tapered $\emptyset$-run $r$ with $r(w)=\emptyset$, for all $w\in W$, which is trivially $1$- and $2$-coherent as well as $1$- and $2$-saturated.
@@
For each $v\in W^{0}$ with $h(v)\geq 0$, we consider the $\Theta^{1}(\Phi)$-quasistate $\boldsymbol{q}^{v}\in\boldsymbol{Q}_{1}(\varphi)$ such that $\mathfrak{M}^{0},v\models(\hat{\boldsymbol{q}}^{v})^{\mathbf{1}}$. Observe that $\boldsymbol{q}^{v}$ may contain formulas $\chi$ with $\textit{md}^{\mathbf{1}}{(\chi)}>h(v)$. Let $\Phi^{v}=\Theta^{\mathbf{1}}(\Phi)\cap\Gamma^{\mathbf{1}}_{h(v)}(\Phi)$ and consider $\boldsymbol{q}^{v}|_{\smash{\Gamma^{\mathbf{1}}_{h(v)}(\Phi)}}$, which is a $\Phi^{v}$-quasistate with $\boldsymbol{q}^{v}|_{\smash{\Gamma^{\mathbf{1}}_{h(v)}(\Phi)}}=\mathfrak{R}^{0}(v)|_{\Theta^{\mathbf{1}}(\Phi)}$. Let $\varphi^{v}$ be its realisability sentence. Since $\textit{adp}^{\mathbf{2}}(\varphi^{v})<k$, we apply the induction hypothesis to the $\mathcal{L}$-formula $\varphi^{v}$ and obtain a tapered 2-cactus $\mathfrak{C}^{v}=((W^{v},R_{1}^{v},R_{2}^{v}),v,\mathfrak{R}^{v})$ for $\Phi^{v}$ (strictly speaking, for $\mathop{\textit{sub}}\varphi^{v}$ but the only difference is that $\mathop{\textit{sub}}\varphi^{v}$ contains Boolean combinations of quantified formulas from $\Phi^{v}$) such that $\boldsymbol{q}^{v}|_{\smash{\Gamma^{\mathbf{1}}_{h(v)}(\Phi)}}=\mathfrak{R}^{v}(v)$. For each $v\in W^{0}$ with $h(v)<0$, we take a tapered 2-cactus $\mathfrak{C}^{v}$ for $\emptyset$. Next, for each $v\in W^{0}\setminus\{w_{0}\}$, we graft the tapered 2-cactus $\mathfrak{C}^{v}$ onto $v$ in the 1-quasimodel $\mathfrak{Q}$ to obtain a tapered 1-cactus $\mathfrak{C}$. A fully grafted tapered 1-cactus can be obtained by grafting, in addition, $\mathfrak{C}^{w_{0}}$ onto $w_{0}$ in $\mathfrak{C}$.
@@ 4 Fusions of One-Variable Logics with Equality
We show that once equality is allowed in one‑variable modal logics, the transfer results established in Section 3 no longer hold. Remarkably, the counterexamples we introduce next arise quite naturally. The propositional modal logic $\mathbf{Diff}$ of the ‘elsewhere’ operator was introduced by Von Wright [32] and is determined by the class $\mathcal{D}$ of all disjoint unions of frames $(W,R)$ with $R=\{(w,v)\in W^{2}\mid w\not=v\}$; see also [8]. We are interested here in the one-variable modal logics $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}$ and $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\text{fin}}$ with $\mathcal{D}_{\text{fin}}$ the class of disjoint unions of finite frames in $\mathcal{D}$. Note that the propositional logic Diff has the ffmp, and so is complete for both $\mathcal{D}$ and $\mathcal{D}_{\text{fin}}$. On the other hand, while one can easily prove the fmp for $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\text{fin}}$, the logic $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}$ does not have the ffmp, and so $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}\subsetneq\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\text{fin}}$ [20]. It has been observed [19, 20] that both logics can be reduced in polytime to the two-variable fragment of first-order logic with counting on arbitrary and finite domains, respectively, and so they are decidable in coNExpTime.
@@
If $e\in E$ is of the form $y=n$, then we take $n$ constants $c_{1}^{e},\ldots,c_{n}^{e}$ interpreted differently and make the set of their interpretations the extension of $P_{y}$:
@@
Figure 2: a) Formula $\textsf{card}(I,P)$. b) Equation $e$ of the form $y=z_{1}\times z_{2}$ with $z_{1}=3$ and $z_{2}=2$.
@@
which implies that equation $e$ holds in the sense that $|P_{y}|_{w}=|P_{z_{1}}|_{w}\times|P_{z_{2}}|_{w}$, for all $w\in W$; see Fig. 2 b), where only worlds in $W$ and their domains are shown.